OpenStudy (anonymous):

If T is the number of teaspoons, then 1 _ 3 T can be used to find the number of tablespoons. Find the tablespoons of spice in each recipe. 5. 3 teaspoons

OpenStudy (anonymous):

you mean 1-3t right?!

OpenStudy (anonymous):

oops yea

OpenStudy (zzr0ck3r):

this makes no sense. f(t) = 1 - 3t is going to give negative values for any amount of t that is positive

OpenStudy (zzr0ck3r):

1 tsp = -2tblspoons....

OpenStudy (anonymous):

honestly i just copy and pasted the question.

OpenStudy (anonymous):

i agree it doesn't make sense

OpenStudy (zzr0ck3r):

well look at it, and then type it. if you cant type it...use the draw function on the bottom left

OpenStudy (anonymous):

can you snapshot it?

OpenStudy (anonymous):

no cause im on the computer but ill try to draw it or type it

OpenStudy (zzr0ck3r):

ok

OpenStudy (anonymous):

here i just attached the file. its question number 5

OpenStudy (zzr0ck3r):

\[f(t) = \frac{1}{3}t\]you need\[f(5.3)=\frac{1}{3}5.3=?\]